SpaceRef

SpaceRef


NOAA GEOS SATOPS Morning Report 3 August 2000

Status Report From: China National Space Administration
Posted: Thursday, August 3, 2000


SATOPS MORNING REPORT - August 3, 2000

GOES

No change in the status of the GOES spacecraft.
 
GOES-10 frame break occurred at 1848Z due equipment problem at Wallops CDA.

GOES-10 rapid scan operations were activated from 2003Z to 0203Z for severe weather over Nevada.

GOES-10 E/W stationkeeping maneuver #13 was successfully executed at 0954Z on August 3, 2000.  Thruster firing and attitude control mode transition activities were nominal.  First post-maneuver image frame started at 1059Z.  Preliminary post-maneuver INR results indicated that the attitude error was small.

POES

No change in the status of the POES spacecraft.

Special NOAA-15 stored AIP testing is being conducted on August 3 to verify and protect quality AMSU datasets independent of the AVHRR/MIRP anomaly.

NOAA-15 HIRS continues to display erratic behavior.  Filter wheel motor current is still within prescribed limit guidelines.

NOAA-15 rev 11546, G5A was noisy from the spacecraft with 90.3% of the data recovered.

DMSP

No change in the status of the DMSP spacecraft.

DMSP operations were nominal over the past 24 hours.
 

Questions or comments to dmiller@nesdis.noaa.gov

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.

SpaceRef Newsletter